In a transportation problem with 4 sources and 6 destinations, how many fixed requirement constraints will be needed?
Answer #1

Number of fixed requirement constraints will be needed = (m + n -1)

Where,

Sources m = 4

Destinations n = 6

Therefore,

Number of fixed requirement constraints will be needed = (4 + 6 -1)

= 10 – 1 = 9

Therefore number of fixed requirement constraints needed is 9